### v1

# GeoGen: Geometry-Aware Generative Modeling via Signed Distance Functions

Abstract
We introduce a new generative approach for synthesizing 3D geometry and images from single-view collections. Most existing approaches predict volumetric density to render multi-view consistent images. By employing volumetric rendering using neural radiance fields, they inherit a key limitation: the generated geometry is noisy and unconstrained, limiting the quality and utility of the output meshes. To address this issue, we propose GeoGen, a new SDF-based 3D generative model trained in an end-to-end manner. Initially, we reinterpret the volumetric density as a Signed Distance Function (SDF). This allows us to introduce useful priors to generate valid meshes. However, those priors prevent the generative model from learning details, limiting the applicability of the method to real-world scenarios. To alleviate that problem, we make the transformation learnable and constrain the rendered depth map to be consistent with the zero-level set of the SDF. Through the lens of adversarial training, we encourage the network to produce higher fidelity details on the output meshes. For evaluation, we introduce a synthetic dataset of human avatars captured from 360-degree camera angles, to overcome the challenges presented by real-world datasets, which often lack 3D consistency and do not cover all camera angles. Our experiments on multiple datasets show that GeoGen produces visually and quantitatively better geometry than the previous generative models based on neural radiance fields.
1 Introduction
†Work conducted during an internship at Microsoft.
The combination of generative models [26, 21, 19, 20] and implicit neural representations [7, 25, 32] has sparked considerable advancements in 3D representation learning [4, 14]. It has powered the synthesis of high-quality, multi-view consistent, images. However, a common pitfall in the pursuit of higher image quality is the sidelining of the quality of the underlying geometry [43].
Recent non-generative efforts, such as NeuS [39], VolSDF [43], and Geo-Neus [11], have made use of the zero-level set of a Signed Distance Function (SDF) to represent the surface of the geometry in a scene via a surface rendering equation, ultimately achieving high-fidelity scene reconstruction. While these models have shown impressive potential, given their non-generative nature, they are only able to reconstruct a scene of interest when multi-view image data is available. This limitation highlights the need for generative models capable of producing high-quality 2D images that are suitable for content creation while ensuring precise geometric synthesis without multi-view data.
Other recent methods such as Ball-GAN [36], and EG3D [5], have combined generative models with Neural Radiance Fields (NeRFs) [26] to yield high quality rendered images. Yet, these approaches often result in noisy meshes that contain geometric artifacts, which emerge due to the properties of NeRFs and their lack of constraints on the geometric reconstructions. Attempts have also been made to harmonize SDFs with generative models as in [31]. However, the generated meshes are often overly smooth, a result of the smoothing prior that encourages the SDF to produce valid values everywhere in 3D space. Additionally, applying this loss can be prohibitive at higher resolutions.
In this work, we address these issues by adding SDF constraints to improve the synthesized geometry of a 3D-aware generative model. Our approach, named GeoGen, employs an SDF depth map consistency loss for enhanced geometric generation. Specifically, we build on EG3D [5] by introducing an SDF representation, instead of a density representation, to encode the geometry. This allows GeoGen to extract mesh surfaces directly from the zero-level set of the SDF [30, 39, 43]. In order to make the SDF representation learning feasible, and to endow it with the ability to model complex and detailed geometry, we also propose an SDF depth map consistency loss. We use a fixed density-to-SDF transformation function to convert the density representation to an SDF. This facilitates generative feature learning by making the learning objective easier to optimize. The SDF also enables the extraction of smooth depth maps that serve as a ‘pseudo’ ground-truth. Our approach uses its own depth prediction in a self-supervised manner to improve the reconstruction. In contrast to commonly used priors, our approach is cheap to compute with only a minor increase in training time.
GeoGen is able to generate detailed meshes from a single input 2D image via inversion [33]. This capability is valuable in applications where the requirement for detailed and realistic meshes is needed. In stark contrast to recent methods like Rodin [40], which required 30 million images during training to create 3D meshes, GeoGen uses a fraction of this number – approximately 50,000 images. Other methods such as PanoHead [1] propose an augmented triplane and separate foreground and background in 2D images with the help of a custom in-house dataset. However, with our proposed architecture, we show that by enforcing our geometric constraints, we are able to reconstruct a detailed 360 geometry, with a reduction in visual artifacts (e.g. the backs of heads) compared to methods such as EG3D [1].
We make the following contributions: (i) We address the problem of 3D synthesis from 2D images by combining a Signed Distance Function (SDF) network with a StyleGAN generative architecture. Our GeoGen model produces more refined geometry predictions compared to conventional neural volume rendering. (ii) We propose an SDF depth map consistency loss that is designed to address geometric inaccuracies from volumetric integration by aligning 3D points with the SDF network’s zero-level set for more precise reconstructions. (iii) We introduce a new dataset of realistic synthetic human heads that contains 360 camera views from multiple synthetic humans. This dataset will be a valuable resource for training and quantitatively evaluating 3D generative models. It can be found on our webpage https://microsoft.github.io/GeoGen.
2 Related work
The landscape of generative modeling has seen a shift in recent years, with techniques drawing on neural implicit representations, such as Generative Adversarial Networks (GANs) [14] and Diffusion models [10, 18, 24, 37] emerging as powerful tools. These techniques blend generative models with neural volume rendering, thereby synthesizing 3D images that capture novel viewpoints from 2D data alone [26]. However, a recurring challenge in this domain has been the reliance on generic density functions to learn the geometry of the images, a factor that often introduces artifacts and results in noisy, low-quality geometric predictions [31]. To mitigate this, prior work has taken advantage of large amounts of multi-view data to constrain the models, thereby yielding more robust geometry [39, 43], but at the expense of not being fully generative.
The emergence of volumetric implicit representations, bolstered by the strengths of Multi-Layer Perceptrons (MLPs) [15] and neural rendering techniques [26], has shown substantial promise in extracting detailed geometry from a 3D scene. This is most apparent in methods such as NeuS [39] and VolSDF [43], which extract high-fidelity surfaces by representing the scene using the Signed Distance Function (SDF) and extracting the surface at the zero level set.
Meanwhile, the broader field of deep learning has seen a surge in novel methods for creating 3D representations from 2D data. One such family of methods is Neural Radiance Fields (NeRFs) [26], which employs a neural network to model the radiance of a 3D scene at any spatial point. The ability of NeRFs to generate high-fidelity 3D models from 2D multi-view supervision, complete with accurate lighting and shading effects, makes them an attractive option for applications requiring realistic 3D representations, such as virtual reality [43].
One set of methods that deserves particular discussion within this landscape is the set of 3D-aware generative models [4, 9, 12, 13, 16, 27, 28, 29, 35]. These methods are specifically designed to generate 3D representations of objects or scenes, utilizing a variety of techniques, including volumetric representations, SDFs, and implicit neural representations. For instance, the Generative Radiance Fields (GRAF) model [34] generates high-resolution 3D shapes with intricate detail, leveraging a neural network to model the radiance and shape of a 3D object. Other notable models include DeepSDF [32], which learns continuous signed distance functions for arbitrary shapes using 3D supervision, and HoloGAN [27], which generates 3D objects by imposing structural constraints in the generative process. Recently, EG3D [5] proposed a triplane representation for volume rendering in generative models, which enables efficient 3D-aware generation. However, extracting high quality 3D meshes is not guaranteed because of its use of a volume density representation. StyleSDF [31], makes use of an SDF representation to directly model geometry, but the extracted surfaces are overly smooth making it challenging to use them in practical applications.
In our investigation of 3D-aware generative models and SDF representations, we identify certain limitations inherent in existing methodologies. One such limitation appears to be a result of the use of the Eikonal loss [43, 39, 11], leading to overly smooth geometry synthesis. Our methodology, building on the foundation laid by EG3D, aims to overcome this by introducing an SDF depth-consistency constraint. This novel constraint is designed to refine geometric surface predictions by leveraging a self-supervised depth prediction mechanism. Unlike previous efforts, such as StyleSDF [31], which merely translates SDF values into density fields, our approach harnesses the full potential of SDF for geometry representation as exemplified by VolSDF [43]. We emphasize that incorporating our SDF representation and its associated constraints does not substantially complicate the training of generative models yet provides enhanced control over geometric surface detail.
Figure 1: GeoGen, our 3D-aware generator, is trained solely from 2D images. Noise sampling is followed by a StyleGan2 generator that produces triplane features similar to EG3D [5]. However, we enhance them with positional info and an SDF network for refined geometry. GeoGen is end-to-end trained with a GAN objective along with our SDF depth consistency loss.
3 Method
Here we present our GeoGen generative approach for enhanced geometric synthesis. We begin by revisiting EG3D [5], an efficient geometry-aware 3D GAN that introduces notation and provides context for our contributions. Then we describe our SDF-based generative model which builds on the EG3D framework.
3.1 Efficient geometry-aware 3D GAN
EG3D [5] is an efficient geometry-aware 3D generative adversarial network. It consists of a StyleGAN2 [20] based feature generator, triplane representation, implicit volume render, and super-resolution module. In order to generate an image, it first samples a random latent noise code and processes the code via a mapping network. The processed code is used to drive the StyleGAN2 generator to produce feature maps which are reshaped to form three feature planes. During the volume rendering, a queried 3D point $\bm{p}$ is projected onto each of the three feature planes, leading to corresponding feature vector $[F_{xy}(\bm{p}),F_{xz}(\bm{p}),F_{yz}(\bm{p})]$. These feature vectors are further processed by a shallow MLP to yield the color and density at the position $\bm{p}$. By the process of volumetric integration, a low-resolution image is generated based on the sampled points along all image rays. Finally, a super-resolution module is used to generate high-resolution output images.
Like EG3D, we also use a triplane representation to efficiently generate images. Different from EG3D, which targets geometry-aware image synthesis, we focus on high-quality geometry synthesis. To this aim, we introduce an SDF-based generative model and present a novel SDF learning strategy.
3.2 SDF-based generative model
Our goal is to develop a model that can learn to generate 3D consistent object-centric images with associated geometry by making use of a collection of posed single-view 2D images at training time. This transformation is achieved by conceptualizing the surface as the zero-level set of a neural implicit signed distance function. To achieve our high-fidelity geometric synthesis, we first introduce our augmented triplane representation. Then, we introduce our SDF-based volume rendering. Finally, we describe an SDF depth-consistency constraint, which is used to enhance SDF learning. Figure 1 displays our overall pipeline.
Augmented triplane representation. Our method augments the original EG3D triplane representation with sampling position $\bm{p}$. According to the sampling position $\bm{p}$, we retrieve the corresponding feature vector $[F_{xy}(\bm{p}),F_{xz}(\bm{p}),F_{yz}(\bm{p})]$ via bilinear interpolation. In addition, the position $\bm{p}$ is processed by a position embedder $PE(\cdot)$ that employs multi-level sine and cosine functions similar to NeRFs [26]:
$PE(a)=[a,\gamma_{0}(a),\gamma_{1}(a),\dots,\gamma_{L-1}(a)],$ (1)
where $\gamma_{k}(a)=[\sin(2^{k}\pi a),\cos(2^{k}\pi a)]$, $L$ is a hyper-parameter that controls the maximum encoded frequency, and $a$ represents each of the three different spatial dimensions of $\bm{p}$. $\bm{p}$ is defined as a vector since it represents the position in 3D space. Each component of $\bm{p}$ (i.e., $p_{x}$, $p_{y}$, $p_{z}$) corresponds to a different spatial dimension.
The function $\gamma_{k}(a)$ is a positional encoding function that takes a scalar value $a$ and returns a 2D vector representation of the sine and cosine of $2^{k}\pi a$. This function is used for positional encoding to capture frequency information up to a maximum frequency defined by the hyper-parameter $L$.
The augmented triplane representation is formed by concatenating the triplane features $F_{xy}(\bm{p})$, $F_{xz}(\bm{p})$, and $F_{yz}(\bm{p})$ with the positional encoding $PE(p_{x})$, $PE(p_{y})$, and $PE(p_{z})$. This augmented representation enables the model to capture high-frequency details by combining the local geometric features with positional encoding information. The absence of the positional encoder destabilizes the training process, often resulting in model collapse (see supplementary material for results).
SDF-based volume rendering. The augmented tri-plane representation is directed to a shallow MLP to learn the SDF value $s(\bm{p})$ and RGB color ${\bf c}(\bm{p})$ for point $\bm{p}$. The SDF value represents the distance to the surface, providing an accurate depiction of its geometry. To convert the SDF value $s(\bm{p})$ into a density field $\sigma$, we follow VolSDF [43] and use the following Laplace transformation:
$\sigma(s(\bm{p}))=\begin{cases}\frac{1}{2\beta}\exp\left(\frac{s(\bm{p})}{\beta}\right)&\text{if }s(\bm{p})\leq 0\\ \frac{1}{\beta}\left(1-\frac{1}{2}\exp\left(-\frac{s(\bm{p})}{\beta}\right)\right)&\text{if }s(\bm{p})>0\end{cases},$ (2)
where $\beta$ is a parameter, which can be fixed or learned. Based on the volumetric integration, the rendered RGB color for a ray ${\bf r}(t)={\bf o}+t{\bf d}$ is calculated as follows:
$C({\bf r})=\sum_{i=1}^{M}T_{i}(1-\exp(-\sigma_{i}\delta_{i})){\bf c}_{i},$ (3)
where ${\bf o}$ is the camera position, ${\bf d}$ is the ray direction, $T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$ and $\delta_{i}=t_{i+1}-t_{i}$ is the distance between adjacent sampled points. For simplicity, we use $\sigma_{i}$ and ${\bf c}_{i}$ to denote $\sigma(s(\bm{p}_{i}))$ and ${\bf c}(\bm{p}_{i})$ respectively, which mean the color and density value at the i-th sampling point $\bm{p}_{i}$ along ray ${\bf r}$. In a similar way, we compute the rendered distance as follows:
$d({\bf r})=\sum_{i=1}^{M}T_{i}(1-\exp(-\sigma_{i}\delta_{i}))t_{i}.$ (4)
SDF depth consistency. It has been shown in Geo-Neus [11] that there can exist a gap between the rendered image and the true surface and it is important to introduce explicit constraints to optimize the SDF network. Therefore, Geo-Neus introduces sparse points and multi-view photometric consistency to achieve this in the multi-view setting when multiple images are available for each object during training. However, these two constraints are obviously not available in our single-view GAN setting. To reduce the geometry bias caused by volumetric integration, the 3D point computed from the rendered distance $d({\bf r})$ in Equation 4 should be located on the zero-level set of the SDF network. Thus, according to the rendered distance $d({\bf r})$, its corresponding 3D point ${\bf p}_{d({\bf r})}$ is computed as:
${\bf p}_{d({\bf r})}={\bf o}+d({\bf r}){\bf d}.$ (5)
Since the above 3D point should be approximately on the geometry surface, the SDF value of this point should be approximately zero. Thus, we define an SDF constraint as:
$\mathcal{L}_{s}=\frac{1}{|\mathcal{R}|}\sum_{\bf r\in\mathcal{R}}|s({\bf p}_{d(\bf r)})|,$ (6)
where $\mathcal{R}$ denotes all rays for the current camera pose. During training we aim to minimize the above loss.
3.3 Training GeoGen
The SDF-based GeoGen model uses dual discrimination during training, evaluating both the neurally rendered low-resolution 2D image and the super-resolved 2D image. The generative model takes only 2D images as input, and the discriminator encourages both the low-resolution and super-resolved synthesized 2D images to match the distribution of real images. This ensures the consistency between the super-resolved images and the neural rendering, facilitating our method to achieve high-quality high-resolution rendering results. In addition, the SDF depth consistency loss is imposed during training to promote geometric consistency. The model can then effectively learn to capture accurate geometry information from the 2D images, leading to more precise and reliable 3D reconstructions. Our overall loss is:
$\mathcal{L}=\mathcal{L}_{dis}+\lambda\mathcal{L}_{s},$ (7)
where $\mathcal{L}_{dis}$ is a GAN loss computed using dual discrimination and $\lambda$ is a weighting applied to the SDF constraint. Empirically we find that directly training our model from scratch is challenging. We suspect that the introduced learnable parameter $\beta$ in Equation 2 prevents the StyleGAN2-based feature generator from learning effective features. In addition, the SDF constraint requires good geometry initialization, which is not possible to obtain in the early phase of training. Therefore, we design a learning strategy to train our model in which the $\beta$ parameter of the Laplace density distribution is fixed to stabilize the early learning of our generative model.
Specifically, the significant part of this training process involves managing the $\beta$ parameter of the Laplace transformation in Equation 2, which directly influences the learning of the SDF network. The $\beta$ parameter remains fixed for the first $N$ iterations to allow the SDF network to focus on learning coarse geometry. This enables the learning of the StyleGAN2-based generator to produce stable view synthesis. After $N$ iterations, we make the $\beta$ a learnable parameter to increase the ability of the model to capture finer-scale surface details. As previously mentioned, the SDF constraint should also be carefully managed. We achieve this by controlling the weight $\lambda$ in Equation 7, where it is initially set to 00, and then increased to $0.1$ after $N$ iterations. As a result, our geometry optimization is conducted in a quasi coarse-to-fine fashion, i.e. $N$ iterations, our Geo-Gen learns coarse geometry and then after this, the SDF constraint can concentrate on surface detail refinement.
4 Synthetic human head dataset
Existing methods typically train their models on high resolution human face datasets such as Flickr-Faces-HQ (FFHQ) [19]. However, FFHQ only contains a limited range of captured viewpoints (i.e. no backs of heads) and has no 3D ground-truth, hence the need for our synthetic dataset. There are other synthetic datasets, such as ShapeNet Cars [6], which have ground-truth 3D meshes but are not realistic looking.
To address this, we created a new dataset of semi-realistic synthetic human heads which is generated based on the work of Wood et al. [41]. Our dataset features images of different synthetic individuals with diverse facial features, body morphologies, clothing, and hair styles. Crucially, unlike FFHQ which primarily captures frontal views, our dataset includes images across the full azimuth range, ensuring comprehensive representation of heads from all sides. This approach not only fills a critical gap in available resources but also shifts the focus towards the quality of the mesh, a vital aspect for advancing the field of 3D generative modeling.
For our dataset, we randomly generate 10 images of 512$\times$512 for each of 19,800 identities, ensuring a comprehensive set of different views, encompassing full azimuthal coverage and utilize multi-view stereo and surface reconstruction techniques to establish pseudo ground-truth meshes. To generate a pseudo ground-truth mesh for quantitative evaluation of 3D reconstruction metrics we use the ACMP multi-view stereo approach from [42] and Poisson surface reconstruction [22] to reconstruct the full head geometry. Example images can be found in Figure 2. A subset of images from our synthetic dataset will be made available upon acceptance.
Figure 2: Examples from our synthetic human dataset. We display rendered images on top and pseudo 3D ground-truth below.
[…]
FFHQ GRAF 79.20 55.00 -
PiGAN 83.00 85.80 0.67
GIRAFFE 31.20 20.10 0.64
HoloGAN 90.90 75.50 -
StyleSDF 11.50 2.65 -
EG3D 4.86 0.0053 0.77
EG3D (rebalanced) 4.70 0.0044 0.79
EG3D∗∗ 5.70 0.0054 0.76
GeoGen 5.40 0.0049 0.75
Synthetic Heads EG3D∗∗ 5.90 0.65 0.69
GeoGen 5.10 0.0038 0.69
ShapeNet Cars GIRAFFE 27.30 1.70 -
Pi-GAN 17.30 0.93 -
EG3D 2.75 0.0054 -
EG3D∗∗ 2.90 0.0043 -
GeoGen 2.50 0.0028 -
Table 1: Comparative analysis of different generative models on FFHQ, our Synthetic Heads, and ShapeNet Cars datasets using standard 2D metrics. Our model surpasses EG3D [5] and other leading models in both FID and ID metrics for the Synthetic Heads and ShapeNet V1 datasets. However, it does not outperform EG3D on the FFHQ dataset, attributed to a lower number of training iterations due to limited computational resources. Additionally, the original number of training epochs for achieving the reported FID results in EG3D is not specified by its authors. GeoGen was not included in training on the FFHQ rebalanced dataset due to its unavailability during the training period. ∗∗ indicates our retraining with far fewer iterations and computation power.
5 Experiments
Here we present qualitative and quantitative results comparing GeoGen to existing methods. For the baseline EG3D model, we retrained it on each of the evaluation datasets so that the training settings were consistent with our approach (e.g. the same number of training epochs). Implementation details are provided in the supplementary material.
EG3D
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
StyleSDF
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GeoGen
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Sampled images and meshes from EG3D, Style SDF, and our GeoGen approach on FFHQ. GeoGen meshes display smoothness, anatomical accuracy, and detailed facial features. In contrast to EG3D and Style SDF, GeoGen synthesizes finer geometric detail.
EG3D
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
StyleSDF
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GeoGen
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Sampled images and meshes from EG3D, StyleSDF, and our GeoGen approach trained on our synthetic human head dataset. GeoGen results in fewer overt visual artefacts and more faithfully captures the backs of objects (e.g. see second last column). While the 2D images from the competing methods look plausible, the underlying 3D mesh is not always consistent.
5.1 Datasets
We perform experiments on Flickr-Faces-HQ (FFHQ) [19], ShapeNet Cars [6], and our synthetic human dataset described previously. Each provide distinct, valuable resources for training and evaluating 3D-aware generative models. The FFHQ dataset consists of high-quality real 2D face images. It contains over 70,000 1024$\times$1024 resolution images. ShapeNet Cars provides images for a variety of car models imaged from different viewpoints. The dataset we used for training contains 2,100 different car instances, each with 20 images from different viewpoints.
Input image Pseudo Ground-truth EG3D GeoGen EG3D Zoom GeoGen Zoom
0
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
90
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
270
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Inversion Results for EG3D and GeoGen Models: The figure presents a comparison at 0, 90, and 270 angles to highlight variations in the reconstruction of facial features by the two models.
Yaw: 0° Yaw: 45° Yaw: 90° Yaw: 135° Yaw: 270°
EG3D
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GeoGen w/o SDF DL
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GeoGen
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Yaw: 0° Yaw: 45° Yaw: 90° Yaw: 135° Yaw: 270°
EG3D
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GeoGen w/o SDF DL
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GeoGen
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Comparison of EG3D and GeoGen, with and without SDF Depth Loss (SDF DL) constraints, showing sampled images from models trained on our synthetic human images. These examples highlight GeoGen’s ability to represent finer geometric details, e.g. the ears have more detail than those generated by EG3D. We also observe a failure for EG3D in the top right, where the back of the head contains facial geometry. More qualitative results highlighting the differences in the use of the SDF depth loss are shown in the supplementary.
5.2 Quantitative results
We adopt the widely used Frechet Inception Distance (FID) [17] and Kernel Inception Distance (KID) [2] metrics to measure the image synthesis quality of our GeoGen approach. We also assess multi-view facial identity consistency (ID) by calculating the mean Arcface [8] cosine similarity score between pairs of views of the same synthesized face rendered from random camera poses. We report the results of our retrained EG3D baseline using the same training conditions and our GeoGen model on the three different datasets in Table 1. Our improved results show that our GeoGen can achieve better image synthesis results on synthetic humans and ShapeNet Cars datasets.
An important feature of our approach is its ability to generate accurate meshes from a single image. However, it is difficult to evaluate the geometric quality of generative models on real images as ground-truth 3D shape information is challenging to obtain. Instead, it is possible to obtain the ground-truth meshes for both synthetic datasets that we use. To evaluate the generated meshes of different methods quantitatively, we leverage the GAN inversion technique PTI [33]. Then, given an image from the test set dataset, we can estimate the corresponding latent code by PTI. With the latent code, we can generate both the synthesized image and mesh. In this way, we can compute a range of 3D evaluation metrics that compare the differences between the synthesized mesh and ground-truth mesh to measure the geometry fidelity. Results are presented in Table 2, where we observe that our GeoGen outperforms EG3D.
Figure 7: Comparison of mesh predictions on ShapeNet Cars. Meshes are obtained by inverting the source image to derive latent codes. EG3D meshes display diminished shape fidelity and surface detail. Using SDF constraints in GeoGen improves detail, evident around car wheels and windows. Results for GeoGen without SDF constraints are also shown for context.
5.3 Qualitative results
Here we present qualitative results where we compare GeoGen to existing methods. In Figures 5 and 7 we compare 2D image synthesis of different methods via GAN inversion. We observe that GeoGen results in outputs that more closely match the input image. In Figure 7 we observe that GeoGen captures details such as the spacing between the car body and wheel and, in some instances, even the handles on the doors of the cars. Finally, in Figures 3 and 4 we display sampled outputs (i.e. not inversions).
ShapeNet Cars
[…]
EG3D 0.31 0.31 0.85 0.44 0.33
GeoGen w/o SDF&Depth Loss 0.27 0.28 0.77 0.42 0.31
GeoGen 0.25 0.27 0.77 0.40 0.29
Synthetic Heads
[…]
EG3D 0.21 0.29 0.65 0.54 0.35
GeoGen w/o SDF& Depth Loss 0.19 0.29 0.59 0.45 0.26
GeoGen 0.17 0.27 0.56 0.43 0.24
Table 2: Comparison of different 3D reconstruction metrics for generative models on ShapeNet Cars and our Synthetic Heads dataset. We report averages for MSE, HD, and MSD metrics. Variations of GeoGen without the SDF and Depth Loss constraints are also shown. Best methods for each dataset are bolded.
6 Discussion
Our evaluation shows the competitive performance of our proposed GeoGen model, both qualitatively and quantitatively. To gain deeper insight into the effectiveness of our approach, we employed a suite of metrics that assess both the 2D and 3D aspects of the images and meshes generated by our model. Two quantitative performance areas are of particular note: the synthesis of high-quality 2D images and precise 3D geometric predictions. Our model competes closely with EG3D [26] in terms of 2D metrics, outperforming both StyleSDF [32] and GRAF [34]. This demonstrates our model’s ability to generate high-fidelity 2D images.
Table 2 showcases a systematic comparison between GeoGen and EG3D, revealing the advantages of incorporating Signed Distance Functions (SDF) and SDF depth constraints during training. The lower Chamfer Distance for GeoGen compared to EG3D for both Cars and synthetic human heads is indicative of a more precise alignment between the reconstructed points and corresponding points in the ground-truth. This highlights an improved precision in point-to-point correspondence which is an essential part of 3D reconstruction. The Earth Mover’s Distance, another vital metric in understanding the geometrical congruence between shapes, is also consistently lower for GeoGen. This indicates that the shapes are more similar, requiring fewer alterations to match the ground-truth, thus showing an underlying efficiency in GeoGen’s modeling approach. Finally, the Mean Surface Distance adds to the evidence of GeoGen’s superiority, as it also yields consistently lower values. The implication here is a closer similarity between the reconstructed and target shapes, providing further evidence for GeoGen’s effectiveness.
The utilization of the SDF in GeoGen ensures better geometric consistency in the reconstruction, as it leverages the implicit representation of the mesh’s surface. GeoGen, with its additional depth constraints, preserves topology and fine details that are often overlooked with conventional generative techniques like EG3D (see Figure 6). It is also noteworthy that these numerical advantages, though significant, do not fully represent the perceptual quality of the reconstructed models. Qualitative evaluations indicate that models generated by GeoGen often appear more realistic and accurate, underscoring GeoGen’s advantage in bridging quantitative performance with perceptual realism.
Limitations. Our GAN-based approach, like others, requires posed images for training. Camera poses can be estimated similar to methods used in FFHQ. While we aim to align the expected depth with the SDF’s zero-level set, extending the SDF consistency loss to other points along the ray could theoretically enhance geometric accuracy. However, this would substantially increase computational load. There are also inherent limitations in learning-based methods, such as potential bias from unrepresentative training data, notably in web-scraped human face images.
7 Conclusion
We presented GeoGen, a novel 3D-aware generative model for synthesizing high-quality 2D images with associated accurate 3D geometry, that is trained from 2D images. GeoGen outperforms established methods on several performance metrics. By harnessing the power of neural implicit representations and neural signed distance functions, we have developed a solution that delivers both quality and versatility in the context of 3D representation learning. In addition, we presented a new synthetic human head dataset for training and quantitatively evaluating 3D generative models. GeoGen moves us closer to the goal of enriching fields such as character animation, gaming, and virtual reality with plausible 3D geometry from single input images. Our results affirm the potential of our approach and its relevance in this rapidly evolving field.